Compare the gravitational force of attraction on a 1.0kg object when it is at the surface of the moon to the gravitational force on the same object when it is at the surface of earth.(mass of the moon =7.35 X10^22kg,radius of the moon = 1.78x10^6m,mass off earth = 6x10^24;radius of earth 6.4x10^6m)


1
Expert's answer
2021-12-02T10:08:44-0500

gravitational force:

"F=\frac{GMm}{R^2}"


for moon:

"F_1=\frac{6.67\cdot10^{-11}\cdot7.35\cdot10^{22}}{(1.78\cdot10^6)^2}=1.55" N


for earth:

"F_2=\frac{6.67\cdot10^{-11}\cdot6\cdot10^{24}}{(6.4\cdot10^6)^2}=9.77" N


"F_2\/F_1=9.77\/1.55=6.3"